OpenStudy (anonymous):

An ancient bone contains 77g of carbon and has a decay rate of 45 counts/min. Given carbon's half-life is 5730 years, how old is it? (Assume that the ratio of isotoopes in the atmosphere is Carbon 14/Carbon 12 = 1.3*10^-12 and has remained constant.)

OpenStudy (anonymous):

We are given \[m = 77 g\] \[R = 45 counts/\min=0.75counts/s\] \[t _{1/2} = 5730 years \approx 1.81 \times 10^{11} s\] The age of a sample is calculated using \[\ln( R/ R _{0}) = -\lambda \times t \][1] R0 and lambda are unknown quantities so let's find them. \[\lambda = \ln2/t _{1/2} \][2] lambda - disintegration constant t1/2 - half-life in seconds NOT in YEARS. So we got it. Next thing to calculate is R 0, which is decay rate at t = 0s or at the very beginning of existence of a sample. \[R _{0}=\lambda \times N _{0}\][3] We don't know N0, the number of Carbon 14 isotope atoms at the moment of time t = 0s \[N _{0}=(initial mass-of-carbon-14)/(mass-of-1-atom-of-carbon-14)\] [4] Let us assume that mass of a sample through ages didn't change, cause the change in mass is negligibly small. Everything whenever created on Earth with carbon contained the carbon 14 isotope in the ratio given, thus \[initial-mass-of-carbon-14 = 77 \times(1-1/(1+1.3 \times 10^{-12}))=1.001 \times 10^{-10}g\] so using [4] you will get N0 then using [3] R0. And now you are ready to find the time of life t using [1] equation. I got 25 560 years.

OpenStudy (anonymous):

I'm pasting again formulas that was trimmed \[N _{0}=InitMassOfCarbon14/MassOfOneAtomOfCarbon14\] [4] \[InitMassOfCarbon14=77 g \times(1-1/(1+1.3 \times 10^{-12}))=1.001 \times 10^{-10} g\]

OpenStudy (anonymous):

@Mythias